# MOREL: Enhancing Adversarial Robustness through Multi-Objective Representation Learning

### v2

@@ 4.3 Evaluation and Analysis of black-box Robustness
CIFAR-10 FGSM PGD-20 PGD-100 $\text{CW}_{\infty}$ Avg-Robust best last best last best last best last best last TRADES 82.57 $83.68$ 83.24 $84.24$ 83.14 $84.01$ 84.40 $85.13$ 83.34 $84.27$ MART $80.23$ 84.31 $81.13$ 84.75 $80.93$ 84.67 $82.33$ 85.86 $81.16$ 84.90 MOREL-T 83.25 83.84 83.98 84.34 83.90 84.22 85.09 85.44 84.06 84.46 MOREL-M $80.63$ $82.42$ $81.33$ $83.19$ $81.05$ $82.97$ $82.47$ $84.32$ $81.37$ $83.23$ CIFAR-100 FGSM PGD-20 PGD-100 $\text{CW}_{\infty}$ Avg-Robust best last best last best last best last best last TRADES 56.43 $56.19$ 56.53 $56.25$ $56.39$ $56.12$ $58.07$ $57.80$ 56.86 $56.59$ MART $54.28$ $55.73$ $54.49$ $56.10$ $54.32$ $55.89$ $56.21$ $58.02$ $54.83$ $56.44$ MOREL-T $55.99$ 56.53 $56.49$ 56.87 56.40 56.66 58.36 58.38 $56.81$ 57.11 MOREL-M 58.82 59.63 59.30 59.92 58.98 59.54 61.22 62.02 59.58 60.28
In addition to white-box attacks, we evaluate the robustness of our models against black-box attacks, where the adversary does not have direct access to the model’s parameters or gradients. Table 4.3 presents the performance of MOREL-T and MOREL-M compared to TRADES and MART on both the CIFAR-10 and CIFAR-100 datasets.
For the CIFAR-10 dataset, MOREL-T demonstrates strong robustness against all attacks, achieving an accuracy of $85.09\%$ against the $\text{CW}_{\infty}$ attack, outperforming both TRADES and MART when considering the best models. While MART shows superior performance for the last models, MOREL-T maintains consistently high robustness across both the best and last models. In the CIFAR-100 dataset, MOREL-M excels in FGSM attacks, achieving the highest accuracy of $58.82\%$ among the best models and an accuracy of $59.63\%$ among the last models, outperforming all other methods. Under PGD-20 and PGD-100 attacks, MOREL-M continues to demonstrate strong robustness, with best accuracies of $59.30\%$ and $58.98\%$, respectively. In the $\text{CW}_{\infty}$ attack scenario, MOREL-T also performs well, achieving $58.36\%$ with its best model and $57.11\%$ with its last model. For Avg-Robust, MOREL-M leads the performance in CIFAR-100 with scores of $56.81\%$ (best) and $57.11\%$ (last), indicating strong overall robustness under black-box attacks.
The results in black-box settings further reinforce the effectiveness of our multi-objective learning framework. Particularly in the challenging CIFAR-100 dataset, MOREL-T and MOREL-M consistently outperform TRADES and MART. This indicates that our method effectively generalizes across attack types.
@@ 5 Conclusion
In this paper, we introduced MOREL, a multi-objective feature representation learning framework aimed at enhancing the adversarial robustness of deep neural networks. MOREL encourages the alignment of natural and adversarial features through the use of cosine similarity and contrastive losses during training, promoting the learning of robust feature representations. Our approach consistently outperformed existing methods that similarly require no architectural changes or test-time data purification, such as TRADES and MART, in terms of robustness against a wide range of adversarial attacks, while maintaining high clean accuracy. Moreover, the ability of our multi-objective optimization approach to generalize across various datasets and attack types, without requiring architectural modifications, makes it a practical and scalable solution for real-world applications. For future work, we plan to investigate the transferability of robust features learned by MOREL across different tasks and domains, which could unlock new possibilities for applying adversarially robust models in areas like transfer learning and domain adaptation. Additionally, we aim to explore grouping techniques to extend MOREL’s application to scenarios with limited labeled data, such as semi-supervised or few-shot learning settings.

### v4

@@ 4.1 Implementation Details
WideResNet34-10 FGSM PGD-20 PGD-100 $\text{CW}_{\infty}$ Avg-Robust best last best last best last best last best last CIFAR-10 TRADES 82.57 $83.68$ 83.24 $84.24$ 83.14 $84.01$ 84.40 $85.13$ 83.34 $84.27$ MOREL($\leftarrow$ TRADES) 83.25 83.84 83.98 84.34 83.90 84.22 85.09 85.44 84.06 84.46 MART $80.23$ 84.31 $81.13$ 84.75 $80.93$ 84.67 $82.33$ 85.86 $81.16$ 84.90 MOREL($\leftarrow$ MART) $80.63$ $82.42$ $81.33$ $83.19$ $81.05$ $82.97$ $82.47$ $84.32$ $81.37$ $83.23$ CIFAR-100 TRADES 56.43 $56.19$ 56.53 $56.25$ $56.39$ $56.12$ $58.07$ $57.80$ 56.86 $56.59$ MOREL($\leftarrow$ TRADES) $55.99$ 56.53 $56.49$ 56.87 56.40 56.66 58.36 58.38 $56.81$ 57.11 MART $54.28$ $55.73$ $54.49$ $56.10$ $54.32$ $55.89$ $56.21$ $58.02$ $54.83$ $56.44$ MOREL($\leftarrow$ MART) 58.82 59.63 59.30 59.92 58.98 59.54 61.22 62.02 59.58 60.28
4.2 White-Box Robustness and Performance Evaluation under AutoAttack
In this section, we evaluate the adversarial robustness of our proposed MOREL method under AutoAttack and white-box (where the adversary has full access to the model’s parameters and gradients) attack scenarios. With the ResNet18 architecture (Table 3.1.2), MOREL($\leftarrow$ TRADES) stands out on all three datasets, achieving superior clean accuracy with its best model. It also consistently outperforms TRADES with both its best and last models against AutoAttack on all three datasets, particularly across every evaluated attack on CIFAR-10. Notably, throughout all the considered datasets, MOREL($\leftarrow$ TRADES), MOREL($\leftarrow$ MART) and MOREL($\leftarrow$ LOAT) each consistently outperform their respective baselines by delivering superior average robust performance along with enhanced clean accuracy. This indicates that our defense framework is effective at preserving natural feature representations while simultaneously enhancing robustness. Besides ResNet18, we also evaluate the performance of TRADES, MART, and their MOREL variants under white-box attacks using the WideResNet34-10 architecture (Table 3.1.2). On CIFAR-10, under the PGD-100 attack MOREL($\leftarrow$ MART) consistently outperforms both MART and TRADES by more than $5\%$ with its last model; moreover under the $\text{CW}_{\infty}$ attack, it maintains its dominance with both the best and last models. As shown in Figure 2 for a ViT model (patch size $4$, input size $32\times 32$) evaluated under AutoAttack, our MOREL variants surpass the GAT [7] defense method on CIFAR-10. GAT converts the original model architecture into a Multi-Task Learning (MTL) architecture and trains it using a multi-objective optimization approach with an auxiliary task.
Figure 2: Comparison of defense methods applied to a ViT model on CIFAR-10 under AutoAttack, including GAT [7] evaluated with three auxiliary tasks. SOO stands for Single Objective Optimization methods.
4.3 Transfer-Based Black-Box Robustness and Performance Evaluation under SquareAttack
In addition to white-box attacks, we evaluate the robustness of our models against black-box attacks, where the adversary does not have direct access to the model’s parameters or gradients. Apart from SquareAttack, we consider transfer-based black-box attacks, where adversarial examples are generated using a surrogate ResNet50 model (trained for $200$ epochs) and then transferred to the target models. The surrogate model is trained on clean images using standard training [29]. Consequently, the same attack techniques used in white-box settings are applicable here. For the ResNet18 architecture (Table 4.1), the MOREL variants demonstrate notable gains in robustness across all datasets, as evidenced by their enhanced Avg-Robust scores. Notably, MOREL($\leftarrow$ TRADES) exhibits robust black-box defense performance relative to other methods across all datasets. In particular, its best model on CIFAR-100 outperforms the three baselines by approximately $2.5\%$ in terms of Avg-Robust score. With WideResNet34-10 (Table 4.1), MOREL($\leftarrow$ TRADES) exhibits superior robust accuracy against most attacks compared to TRADES. On CIFAR-100, MOREL($\leftarrow$ MART)’s performance under the $\text{CW}_{\infty}$ attack is particularly notable, outperforming MART by $4\%-5\%$ in both best and last models. The results in black-box settings further reinforce the effectiveness of our multi-objective learning framework, indicating that our method generalizes well across different attack types.
4.4 Ablation Study
We explore the Pareto front by varying the values of the preference vector $k$ for the MOREL framework (MOREL($\leftarrow$ MART)) with a ResNet18 model trained on the CIFAR-10 dataset. Figure LABEL:fig:main_abla provides a visualization of how the loss terms and performance against PGD-20 evolve as we adjust the values of $k_{1}$ (the weight assigned to the robustness loss $\mathcal{L}_{1}$) and $k_{2}$ (the weight assigned to the accuracy loss $\mathcal{L}_{2}$). As the preference shifts from prioritizing robustness ($k_{1}=0.9$) to accuracy ($k_{1}=0.1$), we observe a clear trade-off between the two objectives (Figure LABEL:fig:abla_a). This behavior clearly illustrates the multi-objective nature of the problem, where optimizing for one objective (accuracy or robustness) leads to a trade-off with the other. Figure LABEL:fig:abla_b shows the relationship between robust accuracy and the values of $k_{1}$. As $k_{1}$ decreases towards 0.1, robust accuracy improves, reaching its peak at $k_{1}=0.3$ This emphasizes the importance of appropriately weighting the robustness loss to improve robustness.
Table 5: Clean and robust accuracy of the model with and without the $M_{\text{e}}$ module (and the associated contrastive loss $\mathcal{L}_{csl}$).
$M_{\text{e}}$ (and $\mathcal{L}_{csl}$) Clean 80.09 $80.00$ PGD-20 50.18 $50.03$ PGD-100 49.01 $48.85$ AutoAttack 45.27 $45.15$
Table 6: Comparison of methods with respect to Intra/Inter Class Distance Ratio and Cosine Similarity, considering the features of all training images and their adversarial examples.
Method $d_{\text{intra}}/d_{\text{inter}}$ ($\downarrow$) CS ($\uparrow$) MOREL ($\leftarrow$ TRADES) 0.90649 0.99942 MOREL ($\leftarrow$ MART) 0.89614 0.99815 MOREL ($\leftarrow$ LOAT) 0.89556 0.99790
Table 4.4 compares the performance of MOREL with and without the $M_{\text{e}}$ module. Since its output is only used for computing the loss $\mathcal{L}_{csl}$ (Eq. 9), removing it corresponds to setting $\alpha=0$ (Eq. 10). The robust accuracy under PGD-20 and AutoAttack is higher when the $M_{\text{e}}$ module is present than when it is removed. Similarly, under PGD-100, the model performs better with the Me module ($49.01\%$) than without it ($48.85\%$). These results suggest that the $M_{\text{e}}$ module and contrastive loss $\mathcal{L}_{csl}$ contribute modestly to improving robustness, even against stronger adversarial attacks. Considering the outputs $T$ and $T^{\prime}$ from the MOREL embedding space (Eq. 6), we report in Table 4.4 the average cosine similarities between the features of all training images and their adversarial examples (PGD-10). In addition, considering the combination of clean and adversarial features, we measure the average pairwise distance between features within the same class (intra-class) and across different classes (inter-class), reporting their ratio. A lower ratio indicates better class separation and tighter intra-class clustering. MOREL($\leftarrow$ LOAT) achieves the best Intra/Inter Distance Ratio $(0.89556)$, while MOREL($\leftarrow$ TRADES) achieves the highest cosine similarity $(0.99942)$. MOREL($\leftarrow$ MART) strikes a balance, achieving a strong ratio $(0.89614)$ and cosine similarity $(0.99815)$.
@@ 5 Conclusion
We introduced MOREL, a multi-objective framework that enhances adversarial robustness by aligning natural and adversarial features with cosine similarity and contrastive losses. MOREL outperforms methods like TRADES, MART, and LOAT in robustness and clean accuracy without needing architectural changes at test-time. Its generalizability across datasets and attack types makes it practical for real-world applications. Future work will extend MOREL to scenarios with limited labeled data.
@@ 6 Acknowledgement
Appendix 0.A More Ablation Studies
0.A.1 Evaluating Multi-Objective Optimization Strategies and Dissecting the Impact of Batch Sizes on Model Robustness
In this section, we perform an ablation study of the Multi-Objective Optimization (MOO) method used in the MOREL framework, specifically comparing Weighted Sum (WS) and Conic Scalarization (CS). As shown in Figure LABEL:fig:moo_abla_a, both methods exhibit a convex Pareto front. However, CS (black line) achieves a better balance of the loss functions. Figure LABEL:fig:moo_abla_b compares the robust accuracy of the models trained using WS (red line) and CS (black line) for different values of $k_{1}$, which weights the robustness objective in the multi-objective optimization process. For both WS and CS, the robust accuracy reaches its peak around $k_{1}=0.3$, Where CS achieves the highest improvement, while WS falls behind. The robust accuracy then declines as $k_{1}$ continues to increase. These results highlight the advantages of Conic Scalarization over the standard Weighted Sum in balancing the competing objectives of learning robust features and making accurate predictions in adversarial training, demonstrating superior empirical performance.
In addition, we analyze the impact of varying batch sizes during training and the presence of the $M_{\text{e}}$ module (with $\mathcal{L}_{csl}$) in the embedding space on the model’s robust accuracy. Figure LABEL:fig:moo_abla_c illustrates the overall robust accuracy under PGD-20 attacks as a function of batch size, with values plotted for batch sizes of $8$, $32$, $128$, and $512$. While larger batch sizes are commonly used in contrastive learning to leverage a diverse set of negative samples, our analysis revealed a different dynamic in MOREL. As the batch size increases, the model’s robustness declines. This trend can be attributed to the differences in training paradigms. In standard contrastive learning [14, 4], training typically involves two distinct steps: first, the encoder is trained to cluster features in the embedding space, and then the classifier is trained on top of the frozen encoder. This separation allows larger batch sizes to enhance feature learning by providing a rich diversity of negative samples, with little interference from downstream classification. In contrast, MOREL considers a simultaneous learning approach, optimizing both feature alignment and classification objectives through multi-objective optimization. As these objectives can sometimes conflict, smaller batch sizes seem to focus the optimization process on a narrower subset of samples, reducing the diversity and complexity of competing gradients in each step. This allows the model to resolve conflicts more effectively, maintaining a better balance between the objectives.
0.A.2 Improving MOREL Performance with Appropriate Choice of Feature Type in $\mathcal{L}_{csl}$
While the primary experiments with MOREL were conducted using the natural features $T$ in the $\mathcal{L}_{csl}$ loss (Eq. 9), we additionally present an ablation study considering the adversarial features $T^{\prime}$ and the combination of both natural and adversarial features ($T\bigoplus T^{\prime}$). Interestingly, using $T^{\prime}$ or $T\bigoplus T^{\prime}$ results in improved robustness. These results highlight the potential of incorporating adversarial features in $\mathcal{L}_{csl}$, either alone or in combination with natural features, to enhance the robustness of the MOREL framework, offering valuable insights for future research and applications.
Figure 7: Performance of MOREL ($\leftarrow$ MART) with different feature types in the $\mathcal{L}_{csl}$ loss (Eq. 9): natural ($T$), adversarial ($T^{\prime}$), and combined ($T\bigoplus T^{\prime}$).